OpenStudy (anonymous):

Evaluate the definite integral of the algebraic function. I checked using my calculator and got 38 instead of 34, which, based on my work, is only the result of plugging in the upper limit to the equation without subtracting the result of plugging in the lower limit.

OpenStudy (zzr0ck3r):

function?

OpenStudy (anonymous):

\[\int\limits_{1}^{3}(3x ^{2}+5x-4)dx\]

OpenStudy (zzr0ck3r):

ok so the integral will be \(x^3+\frac{5}{2}x^2-4x|_1^3\) correct?

OpenStudy (zzr0ck3r):

@christina18 ?

OpenStudy (zzr0ck3r):

can you answer my question?

OpenStudy (anonymous):

That's not how I learned how to evaluate, so I'm not sure.

OpenStudy (zzr0ck3r):

\(\int (3x^2+5x-4)dx = x^3+\frac{5}{2}x^2-4x+c \) so \(\int_1^3 (3x^2+5x-4)dx = 3^3+\frac{5}{2}3^2-4*3-(1^3+\frac{5}{2}1^2-4*1) =\\ 3^3+\frac{5}{2}3^2-4*3-1^3-\frac{5}{2}1^2+4*1=38\)

OpenStudy (zzr0ck3r):

you need to take the integral before you evaluate the bounds

OpenStudy (zzr0ck3r):

what you did was \(f(3)-f(1)\) and \(f(3)-f(1)\ne \int_1^3(3x^3+5x-4)dx\)

OpenStudy (zzr0ck3r):

you need to learn how to integrate...

OpenStudy (anonymous):

Ok thank you. I know how to integrate, I don't know why I skipped it. Thank you again.

OpenStudy (zzr0ck3r):

np